A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?


1
Expert's answer
2021-06-27T16:53:42-0400

the given line is

"-7x+3y=-21.5\\

3y=7x-21.5\\

y=\frac{7}{3}x-\frac{21.5}{3}"


here coefficient of x is 7/3 so the slope of the given line is

"m_1=\frac{7}{3}"


now the line PQ should be either parallel or perpendicular 

let the slope of the line PQ is  m2

so when m1=m2  then both lines are parallel

and when m1.m2=-1  then both lines are perpendicular 


"-1.5x-3.5y=-31.5"

"-3.5y=1.5x-31.5"

"3.5y=-1.5x+31.5"

"y=-\frac{1.5}{3.5}x+\frac{31.5}{3.5}"


"y=-\frac{15}{35}x+\frac{315}{35}"


"y=-\frac{3}{7}x+9"


"so\space here \space m_1\ne m_2"


so both lines are not parallel

"now\space find \space its \space product\\

m_1 \cdot m_2= \frac{7}{3}\left(-\frac{3}{7}\right),\,

{\color{Red} m_1 \cdot m_2= -1}"

therefore both lines are perpendicular 


The equation of the central street PQ is

"{\color{Red} -1.5x-3.5y=-31.5}

."